3. Determine the resultant of the force system iven below by the graphical an analytical methods. 95 N 90 N 201 N 80 0 80 800 50 N 800 60 N 2100M
Answer #1

Solution :

Analytical method:

Sum up the X components and Y components of all the vectors .

X = 201 cos 80 + 0 - 95 cos 80 - 50 - 100 cos 80 + 60 cos 80 = ( 201 - 95 - 100 + 60 ) cos 80 - 50

=   11.46 - 50 = - 38.5 N

Similarly Y = ( 201 + 95 - 100 - 60 ) sin 80 + 90 = 223.9 N

resultant force = sqrt{}( X2 + Y2 ) = sqrt{}(-38.5)2 + ( 223.9)2 = 227.2 N

Direction = tan^-1 (223.9 / -38.5) = 80.2 degrees clockwise (North West ) .

75 9° 80